At 200 °C, the concentrations of nitrogen, hydrogen, and ammonia at equilibrium are

measured and found to be [N2] = 2.12 M, [H2] = 1.75 M, and [NH3] = 84.3 M. Calculate

the equilibrium constant at the temperature given.


1
Expert's answer
2022-04-21T16:48:04-0400

N2 + 3H2 = 2NH3

k = [NH3]2 / ([N2] x [H2]3) = 84.32 / (2.12 x 1.753) = 625.5